# Simplifying FFT-based methods for mechanics with automatic differentiation

### v1

@@ 3.3 Application to multiscale problem
For our final application, we showcase how the ease of computation of homogenized stiffness tensor allows seamless integration of the AD-enhanced FFT-based framework with an additional solver. To this end, we explore a multiscale framework where the macroscale problem is tackled using the Finite Element Method (FEM), while the microscale problem employs the FFT-based approach. This example demonstrates the flexibility of the AD-enhanced FFT-based method by integrating it with a FEM solver.
@@
The macroscale problem involves a beam with dimensions $L\times H$, where the left end is fixed in both the x and y directions. A displacement of $10^{-2}L$ is applied to the right edge along the y-axis (see Figure 4a). At the microscale, we use the same composite RVE as in Section 3.2 but with the St. Venant-Kirchhoff material model to capture the geometric nonlinearity (see Equation 5). A phase contrast of 0.01 is chosen between the soft and hard phases to represent voids, thus modeling a hexagonal-lattice metamaterial. We solve the macroscale problem in reference configuration with $\bm{P}$ and $\bm{F}$ as the stress-strain measures. The weak form at the macroscale thus reads,
@@
$\int_{V}\underbrace{\overline{\mathbb{C}}(\bm{F}):\overline{\bm{F}}}_{\overline{\bm{P}}}:\bm{\nabla}{\overline{\bm{u}}}\cdot{}\mathrm{d}V=0,\quad\text{where}\quad\overline{\mathbb{C}}=\dfrac{\partial\overline{\bm{P}}}{\partial\overline{\bm{F}}}\leavevmode\nobreak\ ,$ (6)
@@
where the homogenized tangent operator $\overline{\mathbb{C}}$ is computed from microscale. We discretize the beam into 16 quadrilateral finite elements (16 Gauss points) as shown in Figure 4b, where each Gauss point is linked to the RVE with 1692 voxels. The deformation gradient $\overline{\bm{F}}$ at each Gauss point within the macroscale domain is transferred as input to the corresponding RVE associated with that specific Gauss point. Similar to the previous example, we determine the homogenized tangent operator $\overline{\mathbb{C}}$ through AD and pass it together with the average first-Piola Kirchhoff stress $\overline{\bm{P}}$ to the FEM solver (see Figure 4a). We note that due to a non-symmetric deformation gradient, one needs to use a full deformation gradient (4 components) [31]. This means that $\overline{\mathbb{C}}$ is a $4\times 4$ matrix. We perform the simulation in one displacement increment under static conditions and use FEniCSx to manage the FEM component [15, 22].
@@
Figure 4: Multiscale simulation of a hyperelastic lattice material. (a) Schematic illustration of the beam at the macroscale and the hexagonal lattice structure at the microscale. (b) The deformed state of the beam for an applied displacement of $10^{-2}L$. The color corresponds to the macroscale displacement along the $x-$axis. The finite element discretization is shown by black lines and two Gauss points are highlighted by colored (red and blue) points. (c) The local shear stress distribution within the RVEs associated with the 2 Gauss points highlighted in (b). (d) Normalized shear stress profile along line highlighted by white dashes in (c).
@@
At the microscale, we apply the Newton-Krylov solver (with an outer tolerance of $10^{-5}$ and an inner tolerance of $10^{-6}$) to accurately capture the nonlinearity within the lattices. At the macroscale, a separate Newton-Raphson solver (with a tolerance of $10^{-8}$) is used to ensure proper convergence. We observe that the chosen contrast of 0.01 between the two phases sufficiently replicates the voids, which is demonstrated by the stresses in the soft phases dropping to 0, as shown by the stress profile across the cross-section in Figure 4c. We report that this simulation takes approx. 20 minutes in total (for roughly 1 million degrees of freedom in total). A large part of the computational time is spent in the CG-solver because of the high contrast in the phase properties. This is a known issue for the Fourier-Galerkin approach [40], which can be improved using pre-conditioned solvers [5]. A traditional FE-FFT framework [12], such as finite-difference-based computation of homogenized stiffness tensor, would require 4 operations per Gauss point, one for each component of non-symmetric deformation gradient, to compute $\overline{\mathbb{C}}$. Conversely, the AD-enhanced multiscale framework is notably advantageous in this scenario, as it demands only one operation per Gauss point for the same.
@@ 4 Discussion & Outlook
The integration of automatic differentiation (AD) with Fast Fourier transform (FFT) techniques can considerably simplify the numerical implementation of complex solid mechanics problems. A key advantage is the ease of deriving stress and tangent operators, thereby minimizing human error and enhancing precision. This work combines automatic differentiation with the Fourier-Galerkin method, whose variational structure facilitates straightforward integration of AD. Here, we investigated the potential of AD in FFT-based methods for mechanics, demonstrating how the automatic derivation of stress and tangent operators enables modular FFT approaches. This strengthens the use of FFT-based methods in more advanced models, particularly in multiscale scenarios where computing homogenized tangent stiffness has traditionally been a significant challenge. While we have emphasized potential applications in Section 3, several other research areas could also greatly benefit from combining AD with FFT-based methods.
@@
A notable application in which we expect AD-based FFT-based methods to be beneficial is computational inelasticity. As shown in Section 3.1, AD greatly simplifies the extraction of the local tangent stiffness for inelastic models and is computationally equivalent to analytical expressions. Therefore, one can leverage AD for complex plasticity models such as multi-yield surface J2 plasticity [41], Simo-plasticity [34] or crystal plasticity [42, 43] for which extracting the tangent stiffness is intricate and often relies on numerical differentiation [31]. AD can also streamline the return-mapping algorithm by boosting the robustness of trial-state projections to the yield surface [35, 34]. Furthermore, multiphysical problems (or coupled problems) offer another promising avenue for the application of AD-enhanced FFT techniques [44, 45, 24]. For example, AD can support the creation of coupled or staggered operators, leading to more precise and efficient solutions for such problems. As discussed in [7], linearizing a thermo-mechanical problem requires four-tangent stiffness matrices; using AD, one can greatly simplify this number to 1 or 2. Additionally, expressing them as operators can significantly reduce the memory footprint. We demonstrated that employing AD-derived operators for the St. Venant-Kirchhoff material model, which has a straightforward analytical form (with only a few mathematical operations) for tangent stiffness, can halve memory usage.
Tackling inverse problems such as topology optimization or PDE-constrained problems can also be simplified with the integration of AD into FFT-based methods. For example, a critical operation in such problems is the calculation of the adjoint for sensitivity analysis [21, 46]. This can be computed as a vector-Jacobian operator, a functionality provided by most auto-differentiation libraries [17]. Furthermore, by broadening AD-FFT techniques to encompass second-order homogenization and strain gradient models, one can boost the precision and predictive power of multiscale simulations, thereby expanding the limits of computational mechanics and materials science. This adaptability highlights the potential of AD in enhancing FFT-based methods for numerous mechanical applications.
@@ 5 Conclusion
In this paper, we investigated the use of automatic differentiation (AD) within the FFT-based framework for solid mechanics. We have shown that AD allows the computation of stresses and tangent operators directly from the energy density functional, which makes FFT-based methods more accessible for complex hyperelastic materials. Furthermore, we have demonstrated that AD allows differentiation of intricate tensor operations, FFT transforms, and Newton-Krylov solvers, which enable the computation of tangent operators for materials where local stiffness is not derivable from strain density, e.g., elastoplastic materials. The incorporation of AD into the FFT-based framework decreases the reliance on explicit expressions of stresses and tangent operators while preserving accuracy and computational efficiency. We further employed the AD-enhanced FFT-based method to simplify the computation of homogenized stiffness tensors, a notable computational bottleneck in multiscale problems. Our findings indicate that applying AD within an FFT-based method is highly beneficial and greatly simplifies the implementation of complex solid mechanics problems, as is also evident by the simple code that we provide at [33]. We believe this research will encourage further studies using automatic differentiation within the FFT-based frameworks and will help in advancing the field of computational solid mechanics.

### v2

@@ 4.3 Application to multiscale problem
In this section, we showcase how the ease of computation of homogenized stiffness tensor allows seamless integration of the AD-enhanced FFT-based framework with an additional solver. To this end, we explore a multiscale framework where the macroscale problem is tackled using the Finite Element Method (FEM), while the microscale problem employs the FFT-based approach. This example demonstrates the flexibility of the AD-enhanced FFT-based method by integrating it with a FEM solver.
@@
The macroscale problem involves a beam with dimensions $L\times H$, where the left end is fixed in both the x and y directions. A displacement of $10^{-2}L$ is applied to the right edge along the y-axis (see Figure 7a). At the microscale, we use the same composite RVE as in Section 4.2 but with the St. Venant-Kirchhoff material model to capture the geometric nonlinearity (see Equation 6). A phase contrast of 0.01 is chosen between the compliant and stiff phases to represent voids, thus modeling a hexagonal-lattice metamaterial. We solve the macroscale problem in reference configuration with $\bm{P}$ and $\bm{F}$ as the stress-strain measures. The weak form at the macroscale thus reads,
@@
$\int_{V}\underbrace{\overline{\mathbb{C}}(\bm{F}):\overline{\bm{F}}}_{\overline{\bm{P}}}:\delta\overline{\bm{F}}\cdot{}\mathrm{d}V=0,\quad\text{where}\quad\overline{\mathbb{C}}=\dfrac{\partial\overline{\bm{P}}}{\partial\overline{\bm{F}}}\leavevmode\nobreak\ ,$ (7)
@@
where the $\delta\overline{\bm{F}}$ is the virtual deformation gradient (or test function) and the $\overline{\mathbb{C}}$ is the homogenized tangent operator computed from microscale. We discretize the beam into 16 quadrilateral finite elements (16 Gauss points) as shown in Figure 7b, where each Gauss point is linked to the RVE with 1692 voxels. The deformation gradient $\overline{\bm{F}}$ at each Gauss point within the macroscale domain is transferred as input to the corresponding RVE associated with that specific Gauss point. Similar to the example in Section 4.2, we determine the homogenized tangent operator $\overline{\mathbb{C}}$ through automatic differentiation and pass it together with the average first-Piola Kirchhoff stress $\overline{\bm{P}}$ to the FEM solver (see Figure 7a). We note that due to a non-symmetric deformation gradient, one needs to use a full deformation gradient (4 components) [33]. This means that $\overline{\mathbb{C}}$ is a $4\times 4$ matrix. We perform the simulation in one displacement increment under static conditions and use FEniCSx to manage the FEM component [17, 24].
@@
Figure 7: Multiscale simulation of a hyperelastic lattice material. (a) Schematic illustration of the beam at the macroscale and the hexagonal lattice structure at the microscale. (b) The deformed state of the beam for an applied displacement of $10^{-2}L$. The color corresponds to the macroscale displacement along the $y-$axis. The finite element discretization is shown by black lines and two Gauss points are highlighted by colored (red and blue) points. (c) The local shear stress distribution within the RVEs associated with the 2 Gauss points highlighted in (b). (d) Normalized shear stress profile along line highlighted by white dashes in (c).
@@
At the microscale, we apply the Newton-Krylov solver (with an outer tolerance of $10^{-5}$ and an inner tolerance of $10^{-6}$) to accurately capture the nonlinearity within the lattices. At the macroscale, a separate Newton-Raphson solver (with a tolerance of $10^{-8}$) is used to ensure proper convergence. We observe that the chosen contrast of 0.01 between the two phases sufficiently replicates the voids, which is demonstrated by the stresses in the compliant phases dropping to 0, as shown by the stress profile across the cross-section in Figure 7c. We report that this simulation takes approx. 20 minutes in total (for roughly 1 million degrees of freedom in total). A large part of the computational time is spent in the CG-solver because of the high contrast in the phase properties. This is a known issue for the Fourier-Galerkin approach [43], which can be improved using pre-conditioned solvers [6]. A traditional FE-FFT framework [14], such as finite-difference-based computation of homogenized stiffness tensor, would require 4 operations per Gauss point, one for each component of non-symmetric deformation gradient, to compute $\overline{\mathbb{C}}$. Conversely, the AD-enhanced multiscale framework is notably advantageous in this scenario, as it demands only one operation per Gauss point for the same.
4.4 Beyond stress and tangent computations
The application of automatic differentiation to compute derivatives extends beyond stress and tangent calculations. Several challenges in solid mechanics require derivatives that are difficult to compute, such as evaluating the sensitivity of a complex computational model to its input parameters. In this section, we illustrate two use cases where automatic differentiation can aid in computing sensitivities of complex computational models such as in uncertainty quantification and topology optimization. Here, we employ the FFT-based method, as presented in earlier sections, as the complex computational model.
[…]
4.4.1 Uncertainty Quantification
We investigate how automatic differentiation can assist in optimizing uncertainty quantification within deterministic processes. Uncertainties may arise from variations in material parameters or inconsistencies in material arrangement. A common method to propagate these uncertainties (i.e., calculate the expected and covariance values of the output) employs Monte Carlo simulations, which, while being effective, are computationally demanding. When the input uncertainties are small, the expected value of the output $\mathbb{E}[y]$ and its covariance $\text{Cov}[y]$ can be approximated using the Taylor series expansion. For example, the second-order approximation for the expected value and the first-order approximation for the variance can be expressed as
$\mathbb{E}[y]\approx f(\mathbb{E}[\bm{x}])+\dfrac{1}{2}C_{ij}\dfrac{\partial^{2}f(\bm{x})}{\partial x_{i}\partial x_{j}}\Big{|}_{\bm{x}=\mathbb{E}[\bm{x}]},\quad\mathrm{Cov}[y]\approx C_{ij}\dfrac{\partial f(\bm{x})}{\partial x_{i}}\Big{|}_{\bm{x}=\mathbb{E}[\bm{x}]}\dfrac{\partial f(\bm{x})}{\partial x_{j}}\Big{|}_{\bm{x}=\mathbb{E}[\bm{x}]},$ (8)
where $C_{ij}$ is covariance matrix of input parameters, $\mathbb{E}[\bm{x}]$ is the expected value of the input variables $\bm{x}$, $\partial f/\partial x_{i}$ and $\partial^{2}f/\partial x_{i}\partial x_{j}$ are the first and the second order derivative of a objective function with respect to inputs $x_{i}$. Typically, the covariance matrix of the input variables is known and therefore the task is to find the derivatives. Here, we employ automatic differentiation to derive the first derivative and the second derivative of the objective function $f$ with respect to its input. To this end, we consider the same two-phase material as in Section 4.1.1. The material is hyperelastic and the expected values of the material parameters and their variance are given in Table 2.
$E_{\text{i}}$ [MPa] $\nu_{\text{i}}$ $E_{\text{m}}$ [MPa] $\nu_{\text{m}}$
$\mathbb{E}$ 5.7 0.386 0.57 0.386
$\sqrt{\text{Cov}}$ 0.01 0.03 0.01 0.03
Table 2: The expected and the standard deviation values of the elasticity modulus and the Poisson ratio for the inclusion ($E_{\text{i}}$ and $\nu_{\text{i}}$) and the matrix ($E_{\text{m}}$ and $\nu_{\text{m}}$).
To simplify the analysis, we consider a function $f$ that accepts material parameters as inputs, determines static equilibrium for a macroscopic deformation gradient $\overline{\bm{F}}=[0,0,0.1]$, and then outputs the total strain energy of the two-phase material. Given that $f$ maps a 4-dimensional space to a scalar space, we use reverse-mode AD to efficiently obtain the first derivative. We differentiate this newly derived function using forward-mode to compute the second-order derivative. Since the function that evaluates the first-order derivative maps a 4-dimensional space to another 4-dimensional space, forward-mode proves to be efficient.
We assume that the covariance $C_{ij}$ between the different parameters is known and is computed by drawing 600 samples randomly from a normal distribution. Using the first-order function, the second-order function and the known covariance matrix, we calculate the expected value of the total strain energy and its variance. We note that we only need 2 runs, first for the first-order derivative and second for second-order derivative. The results thus obtained are shown in Figure 8a & b. For reference, we also determine the expected values and variance through Monte-Carlo simulations using the drawn sample of 600. We observe that the AD-computed values match closely with the Monte-Carlo results (see Figure 8)a & b. Additionally, the total time required for the Monte-Carlo method to achieve a saturated value was approximately 7 minutes, whereas the AD-computed uncertainties took around 20 seconds.
4.4.2 Microstructure optimization
Here, we utilize AD-enhanced FFT-based method for topology optimization and use automatic differentiation to simplify the sensitivity calculation. Topology optimization is frequently utilized to design microstructures with tailored homogenized properties, such as achieving the highest bulk modulus or shear modulus with a specific volume fraction. The microstructure, represented as a density field $\rho(\bm{x})$ varying from 1 (material) to 0 (void), is optimized to achieve the desired properties under specified constraints. The optimization problem thus reads:
$\underset{\rho(\bm{x})}{\text{minimize}}$ $:\leavevmode\nobreak\ f(\rho(\bm{x}),\overline{\bm{\varepsilon}})$
s.t $:\leavevmode\nobreak\ \mathcal{F}^{-1}\{\widehat{\mathbb{G}}(\bm{\xi}):\widehat{\bm{\sigma}}(\bm{\xi})\}=\bm{0}\leavevmode\nobreak\ ,$
[…]
: $:\leavevmode\nobreak\ 0\leq\rho(\bm{x})\leq 1,\forall\bm{x}\in\Omega$
where $f$ is the objective function to be minimized. The first condition ensures static equilibrium. The second condition ensures the imposed volumetric constraint and the last condition ensures that the density field $\rho(\bm{x})$ remains within physical limits i.e. $\rho(\bm{x})\in[0,1]$. Irrespective of the solution strategy (optimality criteria [44], MMA [45] or Lagrange multipliers) used for the aforementioned optimization problem, information about the gradients of both the objective function and the constraints with respect to the input is required. Recently, automatic differentiation has been adopted to simplify the computation of these gradients [46]. Building on this concept, here we use automatic differentiation to determine both homogenized properties (i.e. homogenized shear modulus $\overline{\mu}$) and their sensitivities to the density field within a FFT framework.
Similar to [47, 48, 46], we consider the classic microstructure optimization problem of maximizing the homogenized shear modulus $\overline{\mu}$ of a microstructure (i.e. objective function $f(\rho(\bm{x}))=-\overline{\mu}$) with a given volume fraction. We examine a square domain under macroscopic strain $\overline{\bm{\varepsilon}}$. The material behaves as linear elastic with a Young’s modulus $E=1$ MPa and a Poisson ratio $\nu=0.3$. Additionally, we enforce a volumetric constraint of 0.3. We follow the optimality criteria method [44] to solve the above constrained optimization problem. This method requires the sensitivity of the objective function $\partial f/\partial\rho$ to update the density field $\rho$ in successive iterations. Since the function $f$ maps a high-dimensional space ($\rho$) to a scalar value ($\overline{\mu}$), we use reverse-mode AD to compute the sensitivities. For each update of the density field, we use the AD-enhanced FFT framework to calculate the homogenized stiffness tensor $\overline{\mathbb{C}}$ of the microstructure for a macroscopic strain of $\overline{\bm{\varepsilon}}=[0,0,1.]$ under plane-stress conditions, as discussed in Section 4.2.
The optimized microstructure (see Figure 8b) and the converged shear modulus $\overline{\mu}\approx 0.075$ match well the results from the literature [49, 47, 46]. Thus, integrating AD-enhanced computation of homogenized properties with AD-derived sensitivities streamlines the implementation of optimization problems. This approach allows for the simultaneous calculation of all homogenized components, differing from previous studies that required solving the static equilibrium for three separate test strains [50]. Furthermore, integrating simultaneous homogenization of the microstructure and computation of sensitivities within an FFT-based framework facilitates the optimization of computationally intensive cases.
@@ 5 Discussion & Outlook
The integration of automatic differentiation (AD) with Fast-Fourier Transform (FFT) techniques can considerably simplify the numerical implementation of complex solid mechanics problems. A key advantage is the ease of deriving stress and tangent operators, thereby minimizing human error and enhancing precision. This work combines automatic differentiation with the Fourier-Galerkin method, whose variational structure facilitates straightforward integration of automatic differentiation. Here, we investigated the potential of automatic differentiation in FFT-based methods for solid mechanics, demonstrating how the automatic derivation of stress and tangent operators enables modular FFT approaches. This strengthens the use of FFT-based methods in more advanced models, particularly in multiscale scenarios where computing homogenized tangent stiffness has traditionally been a significant challenge. While we have emphasized potential applications in Section 4, several other research areas could also greatly benefit from combining automatic differentiation with FFT-based methods.
@@
A notable application in which we expect AD-based FFT-based methods to be beneficial is computational inelasticity. As shown in Section 4.1, automatic differentiation greatly simplifies the extraction of the local tangent stiffness for inelastic models and is computationally equivalent to analytical expressions. Therefore, one can leverage automatic differentiation for complex plasticity models such as multi-yield surface $J_{2}$-plasticity [51], Simo-plasticity [36] or crystal plasticity [52, 53] for which extracting the tangent stiffness is intricate and often relies on numerical differentiation [33]. Automatic differentiation can also streamline the return-mapping algorithm by boosting the robustness of trial-state projections to the yield surface [37, 36]. Furthermore, multiphysical problems (or coupled problems) offer another promising avenue for the application of AD-enhanced FFT techniques [27, 54, 55]. For example, automatic differentiation can support the creation of coupled or staggered operators, leading to more precise and efficient solutions for such problems. As discussed in [8], linearizing a thermo-mechanical problem requires four-tangent stiffness matrices; using automatic differentiation, one can greatly simplify this number to 1 or 2. Additionally, expressing them as operators can significantly reduce the memory footprint. We demonstrated that employing AD-derived operators for the St. Venant-Kirchhoff material model, which has a straightforward analytical form (with only a few mathematical operations) for tangent stiffness, can halve memory usage.
We showed that automatic differentiation capabilities go beyond just computing stress and tangent values in an FFT-based framework. In Section 4.4, we utilized an AD-enhanced FFT-based method for topology optimization, where an accurate and efficient computation of sensitivity is necessary for a correct and fast optimization. Automatic differentiation simplifies the calculation of the sensitivity by directly differentiating the objective function with respect to the input parameters, thus eliminating the need for adjoint calculations [56]. Likewise, the high computational cost associated with uncertainty quantification, often due to the employed Monte Carlo approach, can be significantly reduced by using higher-order approximations to calculate expected values and variances of an output parameter. Automatic differentiation enables computing higher-order derivatives in a single simulation, thereby significantly reducing the computational cost. Furthermore, by broadening AD-FFT techniques to encompass second-order homogenization and strain gradient models, one can boost the precision and predictive power of multiscale simulations, thereby expanding the limits of computational mechanics and materials science. This adaptability highlights the potential of automatic differentiation in enhancing FFT-based methods for numerous mechanical applications.
The biggest challenge in adding AD techniques to existing codes/libraries is selecting the appropriate AD framework. Many AD frameworks are language-specific, so it is often easier to do so if the code is already in that language. For example, transforming existing Python-based code to work with differentiable frameworks such as JAX is generally straightforward (as is evident in the code provided at [35]). JAX’s functional programming makes it easy to fit into scientific workflows with minimal changes. In contrast, modifying C+++++ + code to function with AD frameworks such as ADOL-C [57], autodiff [58] or Sacado [59] typically requires more changes to handle automatic differentiation structures. However, these C+++++ + libraries, built for performance, excel in tasks with high performance requirements, unlike Python-based framework such as JAX or PyTorch, which incur a slight computational overhead. Although frameworks such as ADOL-C and Sacado enable efficient automatic differentiation for basic arithmetic and algebra operations, allowing the derivation of stresses or tangent moduli from energy density functions, they lack, to the best of authors’ knowledge, native support for FFT differentiation. Therefore, enabling automatic differentiation for FFT functions in these frameworks requires either custom adjoint routines or defining the chain rule for FFT operations, both of which are often complex and time-consuming. In contrast, JAX and PyTorch natively support automatic differentiation through FFTs, as shown in this paper, making them suitable for FFT-based methods. Furthermore, JAX compensates for AD overhead with its just-in-time compilation and advanced parallelization capabilities (CPUs and GPUs). Recently, frameworks such as Enzyme-AD [60, 61], an experimental differentiable library that works across languages, enable automatic differentiation without altering the original code or libraries. This offers both flexibility, differentiation of FFT functions and high performance and, therefore, could be a great balance between ease of integration and computational efficiency.
Another important factor in selecting an AD framework is the range of differentiation techniques it offers, such as forward-mode and reverse-mode. As discussed in Section 2, both methods are crucial and can provide considerable efficiency depending on the function being differentiated. Although PyTorch provides AD support, mainly in reverse-mode, it does not offer as extensive support for both modes as JAX does (forward-mode, jvp function and linearize function). Therefore, the choice of a suitable framework for integrating automatic differentiation into existing code largely hinges on the user’s specific requirements and preferences such as integration ease, flexibility, computing efficiency and support for parallelization. In this research, we opted for JAX to augment FFT-based methods with automatic differentiation. Nonetheless, it is feasible to employ any available AD framework, as the principles addressed in this paper remain applicable.
Finally, in this paper, we employed the Fourier-Galerkin method for the FFT approach; however, the AD technique, as demonstrated here, is applicable to other FFT-based methods [1, 62, 63] as well.
@@ 6 Conclusion
In this paper, we investigated the use of automatic differentiation (AD) within the FFT-based framework for solid mechanics. We have shown that automatic differentiation allows the computation of stresses and tangent operators directly from the energy density functional, which makes FFT-based methods more accessible for complex hyperelastic materials. Furthermore, we have demonstrated that automatic differentiation allows differentiation of intricate tensor operations, FFT transforms, and Newton-Krylov solvers, which enable the computation of tangent operators for materials where local stiffness is not derivable from strain density, e.g., elastoplastic materials. The incorporation of automatic differentiation into the FFT-based framework decreases the reliance on explicit expressions of stresses and tangent operators while preserving accuracy and computational efficiency. We further employed the AD-enhanced FFT-based method to simplify the computation of homogenized stiffness tensors, a notable computational bottleneck in multiscale problems. We show that the AD-enhanced FFT-based approach is applicable beyond stress and tangent calculations, extending to problems like uncertainty quantification and topology optimization by simplifying sensitivity computations. Our findings indicate that applying automatic differentiation within an FFT-based method is highly beneficial and greatly simplifies the implementation of complex solid mechanics problems, as is also evident by the simple code that we provide at [35]. We believe this research will encourage further studies using automatic differentiation within the FFT-based frameworks and will help in advancing the field of computational solid mechanics.